OpenStudy (anonymous):

A nuclear reaction releases 9.00 × 10^13 J of energy. What was the change in mass? 5.0 × 10^-4 kg 1.0 × 10^-3 kg 5.0 × 10^-3 kg 5.0 × 10^-2 kg

OpenStudy (irishboy123):

Here it comes! \(E = m \ c^2\) \(\Delta m = \frac{ E}{c^2}\)

OpenStudy (anonymous):

so E is 9.00 × 10^13, what is c?

OpenStudy (anonymous):

oh nvm is 3.0 x 10^8

OpenStudy (anonymous):

so .001 = 1.0 × 10^-3?

OpenStudy (irishboy123):

looks good